[ how many liters of a 40% alcohol solution must be mixed with 60 liters of a 70% solution to get a 50% solution? ]

Answer by TimothyLamb(4379)

40w + 70(60) = 50V
V = w + 60
---
40w + 70(60) = 50V
40w + 4200 = 50V
40w + 4200 = 50(w + 60)
40w + 4200 = 50w + 3000
10w = 1200
w = 120
---
answer:
need 120 liters of the 40% solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php